The base diameter and altitude of a right circular cone are observed at a certain instant to be 10 and 20 inches, respectively. If the lateral area is constant and the base diameter is increasing at a rate of 1 inch per minute, find the rate at which the altitude is decreasing.


1
Expert's answer
2021-01-25T00:17:07-0500

Given: d = 10, h = 20, S - is constant.

x = 1 (inch/minute) - the rate at which the diameter is increasing.

Find: "y" (inch/minute) - the rate at which the altitude is decreasing.

S0 = d * "\frac{h}{2}"

S0 = 10 * "\frac{20}{2}" = 100.

S1 - the lateral area in one minute.

d1 = d + x

d1 = 10 + 1

d1 = 11.

"h"1 = "h" - "y"

"h"1 = 20 - "y".

S1 = d1 * "\frac{h1}{2}"

S1 = 11 * ("\frac{20 - y}{2}"),


By condition, S1 = S0.

11 * ("\frac{20 - y}{2}") = 100

11 * (20 - "y") = 100 * 2.

20 - "y" = 200 / 11.

"-y" = 18"\frac{2}{11}" - 20.

"-y" = "-1\frac{9}{11}"

"y" = "1\frac{9}{11}".


Answer: "1\frac{9}{11}" inch per minute.